2.1 Consider a closed economy that is described by the following model:



C = 280m + 0.72Y



Where:



C = Consumption Y = Income



I = 150m I = Investment



G = 300m G = Government spending



T = 22% t = Tax rate.




2.1.1 Calculate the multiplier. (3)



2.1.2 Calculate the total autonomous spending. (2)

1
Expert's answer
2022-03-20T19:07:49-0400

2.1.1 Calculate the multiplier. (3)

Solution.

"\alpha" = 1/1(1-C)

C(CMPC)= "\Delta" C/"\Delta" Y = 0.72 =C

"\alpha" = 1/(1-0.72)

= 3.5714

2.1.2 Calculate the total autonomous spending. (2)

Solution

"\Alpha\omicron" = C + I + G

= 280m + 150m + 300m

= 730m